### v1

# Look as You Leap: Planning Simultaneous Motion and Perception for High-d o f Robots

Abstract
In this work, we address the problem of planning robot motions for a high-degree-of-freedom (d o f) robot that effectively achieves a given perception task while the robot and the perception target move in a dynamic environment. Achieving navigation and perception tasks simultaneously is challenging, as these objectives often impose conflicting requirements. Existing methods that compute motion under perception constraints fail to account for obstacles, are designed for low-d o f robots, or rely on simplified models of perception. Furthermore, in dynamic real-world environments, robots must replan and react quickly to changes and directly evaluating the quality of perception (e.g., object detection confidence) is often expensive or infeasible at runtime. This problem is especially important in human-centered environments such as homes and hospitals, where effective perception is essential for safe and reliable operation. To address these challenges, we propose a GPU-parallelized perception-score-guided probabilistic roadmap planner with a neural surrogate model (ps-prm). The planner explicitly incorporates the estimated quality of a perception task into motion planning for high-d o f robots. Our method uses a learned model to approximate perception scores and leverages GPU parallelism to enable efficient online replanning in dynamic settings. We demonstrate that our planner, evaluated on high-d o f robots, outperforms baseline methods in both static and dynamic environments in both simulation and real-robot experiments.
I Introduction
Achieving mobile manipulation (e.g., navigating to a table to grasp a tool) alongside perception tasks (e.g., tracking an object or a human) remains highly challenging. In this work, we address the problem of planning robot motions for a high-degree-of-freedom (d o f) robot that effectively achieves a given perception task while the robot and the perception target move in a dynamic environment. A fundamental challenge emerges: the objectives of motion and perception often conflict—perception tasks impose non-trivial constraints on robot motion, as performance varies with object shape, occlusion, distance, and viewing angle [1, 2, 3]. For example, detecting human faces often requires maintaining a frontal view, while reliable object tracking typically benefits from close proximity. At the same time, motion planning for high-degree-of-freedom (d o f) robots must follow kinematic constraints, avoid obstacles and nearby humans, and account for the complex, nonlinear relationship between the robot’s configuration and the camera pose.
As an example, consider the situation in Fig. 1. A mobile robot with wrist-mounted camera must navigate a cluttered environment while observing a monitor at the end of the hallway. The robot must navigate through clutter with a differential drive base while simultaneously maintaining tracking performance, i.e., it avoids occlusions and attains high tracking confidence.
Figure 1: The robot—a UR5 mounted on a differential drive base—navigates a cluttered environment while maintaining high perception performance. In this example, the robot must continuously track a monitor located down the perpendicular hallway using its wrist-mounted camera. Top: Illustration of the robot’s trajectory—fading from right to left—as it moves while maintaining a high detection rate of the monitor. Bottom: Camera views corresponding to each robot configuration shown in the top figure.
These challenges are further compounded in dynamic environments, where objects and people move. In such settings, robots must continuously adapt to changes while preserving high perception quality [4, 5, 6, 7]. For example, a key motivation of this work comes from human-centered environments such as nurse training settings [8, 9], where robotic tutors are used to support skill acquisition. In such a system [10], the robot must continuously monitor a trainee to detect compliance with sterile techniques and provide timely feedback while simultaneously navigating the environment. Fig. 7 and Fig. 10 illustrate this challenge, where the robot must continuously track nurses’ faces while navigating through a cluttered clinical environment. However, directly evaluating perception performance (e.g., detection confidence) in such settings can be prohibitively expensive or even impossible. This motivates the need for efficient mechanisms to approximate perception quality and support fast, responsive replanning in high-dimensional spaces.
How do existing planning methods tackle planning for perception in high d o f setting? Existing planning methods that consider both motion and perception have several limitations. Informative path planners [11, 12, 13, 14] aim to find paths that maximize information gain while respecting constraints like path length or budget, but they often become computationally expensive, especially when the problem size grows. Perception-aware path planners are typically designed for low-d o f systems [15, 16, 17, 18, 19, 20] or rely on overly simplified perception models [21], such as keeping a point of interest centered in the field of view. Active perception approaches [22, 23, 24, 25, 26, 27, 28, 29, 30] focus on selecting discrete viewpoints to maximize scene understanding, rather than ensuring consistent perception quality along a motion trajectory. Moreover, most of these approaches are not designed to operate in dynamic environments where objects or humans may move and where continuous adaptation is needed. As a result, they lack the responsiveness required for real-time perception-aware planning in practical settings.
In this paper, we propose Perception-Score-guided Probabilistic Roadmap Planning (PS-PRM)—a GPU-parallelized, roadmap-based planner that explicitly incorporates the estimated quality of a given perception task (e.g., object detection confidence) into motion planning for high-degree-of-freedom (DOF) robots. The perception score is predicted by a neural surrogate model trained to approximate perception scores for specific tasks. ps-prm jointly optimizes for both motion and perception by associating each configuration with an estimated perception score and using these scores to guide roadmap construction and path selection. To efficiently approximate perception quality during planning, we introduce a neural surrogate model trained on data from various objects and human targets. To account for occlusions in cluttered environments, we incorporate a ray-casting-based pipeline that dynamically adjusts perception estimates. Finally, to enable fast replanning in dynamic environments, we develop a GPU-parallelized framework that executes collision checking, forward kinematics (FK), and perception evaluation in batch. These components allow our method to scale to high-dimensional planning problems while maintaining robust perception performance in both static and dynamic scenarios.
In this work, we make the following key contributions:
  • A sampling-based roadmap planner, ps-prm, that integrates perception quality into motion planning for high-d o f robots.
  • A perception score estimation pipeline that uses a neural surrogate model and ray casting to efficiently approximate detection quality while accounting for occlusions in cluttered scenes.
  • A batch-processing infrastructure that accelerates collision checking, forward kinematics, and perception evaluation using GPU parallelism, enabling real-time replanning.
  • Extensive simulation and real-robot experiments on three different high-d o f robots, demonstrating that ps-prm consistently improves both perception performance and planning efficiency compared to baseline methods.
II Problem Statement
We consider a robot with $k$ controllable joints, such as a mobile robot with an onboard camera with limited field of view. The robot is described by its configuration space $\mbox{$\mathcal{Q}$}\subseteq\mathbb{R}^{k}$. The robot is located in a workspace $\mbox{$\mathcal{W}$}\subseteq\mathbb{R}^{3}$ that contains objects and people of interest that may need to be monitored, as well as obstacles to be avoided. The robot has access to the pose and geometry of all obstacles in $\mathcal{W}$ and the positions of the centroids of objects of interest. As the robot moves, the objects or humans of interest may also move, and we assume the robot has real-time access to their poses. The subset of all robot configurations that are collision-free is denoted by $\mbox{$\mathcal{Q}_{free}$}\subseteq\mbox{$\mathcal{Q}$}$. The motion planning problem is to find a path $\sigma:[0,1]\rightarrow\mbox{$\mathcal{Q}_{free}$}$ that assigns a collision-free configuration to each input parameter, starting from $\sigma(0)=\mbox{$q_{start}$}$ and ending in a goal region $\sigma(1)\in\mbox{$\mathcal{Q}_{goal}$}$.
We denote the set of all paths in $\mathcal{Q}$ as $\Sigma_{\mathcal{Q}}$, and all collision-free paths as $\Sigma_{\mathcal{Q}_{\mathrm{free}}}$. We additionally define a motion cost function $c_{m}:\Sigma_{\mathcal{Q}}\rightarrow\mathbb{R}^{+}$ and a perception cost function $c_{p}:\Sigma_{\mathcal{Q}}\rightarrow\mathbb{R}^{+}$. The objective of this work is to find a collision-free path $\sigma^{*}:[0,1]\rightarrow\mathcal{Q}_{\text{free}}$, given a path planning problem $(\mathcal{Q}_{\text{free}},q_{\text{start}},q_{\text{goal}})$, that minimizes a composite cost function $c(\sigma)=c_{m}(\sigma)+\alpha\cdot c_{p}(\sigma)$, where the trade-off between motion and perception cost is governed by a scalar weight $\alpha\geq 0$. The optimal path is given by $\sigma^{*}=\arg\min_{\pi\in\Sigma_{\mathcal{Q}_{\text{free}}}}\left[c_{m}(\sigma)+\alpha\cdot c_{p}(\sigma)\right]$, if such a path exists.
In our problem setting, $c_{m}(\sigma)$ represents the exact or approximate arc length of a path $\sigma$ and $c_{p}(\sigma)$ quantifies the performance of a perception task along $\sigma$. For example, in an object tracking task, the perception cost may correspond to the average tracking confidence score obtained as the robot moves along the path. Our problem formulation seeks to minimize a weighted sum of motion and perception costs, aiming to find a path that is both short and achieves good perception performance.
III Related Work
III-A Informative Path Planning
There have been many works on the general problem of informative or information-rich path planning [31, 32, 33]. The general problem of informative path planning is to maximize some metric of information quality (e.g., probability of locating a target, accuracy of the inspection, or quality of the survey) while satisfying constraints on energy, time or perception. The informative motion planning problem of maximizing information gathered subject to a constraint is particularly challenging because it typically requires searching over a large and complex space of possible trajectories. Such problems have been shown to be NP-hard [34], or even PSPACE-hard [35], depending on the form of the objective function and the space of possible trajectories. Several general solvers for robotic information gathering problems utilize combinatorial optimization techniques to search over a discrete grid. The recursive greedy algorithm [34] is one example that achieves bounded performance for submodular objective functions. Branch and bound techniques have also been proposed that only require the objective function to be monotonic [36]. Another approach is to utilize a finite-horizon solver that solves the problem for only a portion of the budget at a time [37]. However, the resulting solvers typically require computation exponential in the size of the problem instance due to the large blowup in the search space. Thus, they are mostly shown with 2D environments and with very few obstacles.
Another related line of works is the planning under uncertainty or planning in information spaces, which focuses on selection of trajectories that minimize the localization uncertainty. This problem has generally been solved with Partially Observable Markov Decision Processes (POMDPs) or through graph-search in the belief space [38]. While these approaches are well-established, in general their computational complexity grows exponentially in the number of possible actions and observations.
To overcome these issues, the use of sampling-based motion planning algorithms, such as the RRT [39] and the probabilistic roadmap (PRM) [40], has increased enormously in past years. Such algorithms have the advantage of quickly exploring a space of interest to achieve a feasible solution. Variants of sampling-based motion planning have also been proposed to maximize some metric of information quality. The prior algorithms have been shown to solve related problems for certain objective functions, including information-rich RRTs [41] using Fisher information matrices and belief-space RRTs [42]. While information-rich sampling based motion planners mentioned above like belief-space RRTs are well-suited for energy minimization tasks, in this work, we specifically focus on selecting trajectories that maximize the perception task performance without considering robot control efforts. Furthermore, these prior algorithms often focus on 2D environments [41, 42, 32, 33] and their information metrics do not apply to the perception task we consider in this work where the approximate analytical model suffers.
III-B Perception-aware Path Planning
Perception-aware path planning has been studied extensively for Unmanned Aerial Vehicles (UAVs) and Micro Air Vehicles (MAVs) because a perception-aware component is especially important when flying in dynamic environments. Furthermore, a consistent detection of the moving obstacles is necessary to obtain a good estimate of their locations and prediction of their future trajectories.
Specifically, our work relates to recent work on optimizing motion and perception tasks for UAVs and MAVs in applications such as quadrotor flying [15, 16, 17, 18, 19, 20], under dynamic environments [11, 12, 13, 14] and for scenes with semantic classes [21]. In most cases, the perception task aims at maintaining visibility of landmarks and informative regions for visual odometry [21], reducing the velocity of the projection of points of interest on the camera plane to avoid blurred motion [15, 43] or guaranteeing the visibility of obstacles for safety [44, 14, 11]. Further relevant work addresses the problem of having a UAV record or chase a target [45, 46, 47, 48] For example, [47] focused on tracking a moving target with a downward-facing camera, while [45] proposed a way to follow a moving target while avoiding other static obstacles in the environment.
For UAVs, most methods formulate the problem as optimization using Model Predictive Control (MPC) [15, 44, 11, 14, 19] or Reinforcement Learning (RL) [12, 17, 18, 49]. MPC-based methods jointly optimize perception and motion for agile UAV navigation using numerical optimization, without obstacles [15] or for static obstacles and occlusions [44]. For example, the work [21] steer the UAV towards informative areas using a path planner while keeping landmarks in the camera field of view through optimization. RL-based methods create policies that directly map from sensor measurements to controls while keeping visibility of regions of interest. The policy learner has access to privileged information during training, e.g., by using MPC [50], a planner [17] or a state-based trained expert [18]. Another line of work focuses on using the UVA or MVA for autonomous aerial cinematography or low altitude traffic surveillance that can reason about both geometry and scene context in real-time [51, 52].
Unlike these works, we consider high-d o f systems—e.g., mobile robots with an onboard camera with limited field-of-view controlled by multiple joints—with kinematic constraints on environments with many obstacles, where optimization is likely to get stuck in a local minima. We focus on simultaneous motion and perception high-d o f robots such as mobile manipulators instead of UAVs or MAVs. Furthermore, unlike most of the perception-aware planning works that aim at maintaining visibility of landmarks and informative regions for visual odometry, we address more complex perception tasks which are significantly more computationally expensive to evaluate.
III-C Active Perception
When perception is incorporated into the path planning process, the problem of selecting optimal viewpoints to maximize the performance of a given task is referred to as active perception or view planning [22, 23, 24]. One of the goals of active perception is active localization, which seeks to compute control actions and trajectories that minimize the pose estimation uncertainty. Most active localization works have been in the context of simultaneous localization and mapping (SLAM) or exploration [25, 30]. Depending on the sensor used, they can be classified into range-based [53, 54, 55] or vision-based [26, 27, 28, 29, 30]. Active perception considers applications such as object [56] and scene [57] reconstruction, object recognition [58], pose estimation [59], and active mapping [25], with single robot or multiple robots [60].
Viewpoint planning or next best view planning (NBV) is a subset of active perception where the sensor pose or a sequence of poses is planned to maximize the information gain, i.e., minimize the entropy or uncertainty about the state of the environment or target objects, subject to constraints such as obstacle avoidance and movement cost. The challenge of next best view planning is proposing and selecting views from which a scene can be efficiently observed. The quality of an observation can be quantified by its accuracy (i.e., how closely the captured data resembles the actual scene) and completeness (i.e., what proportion of the scene has been observed). Observation accuracy primarily depends on the sensor capabilities but can be improved by considering the scene texture and geometry. The completeness of an observation is determined by the coverage obtained from captured views. The efficiency of a NBV approach is quantified by the time and travel distance required to obtain an observation. For detail oriented tasks, especially at the object level, such as active recognition, pose estimation [61], mapping or reconstruction [62, 63], manipulators or mobile manipulators are typically used with attention-driven next best view planning.
Similar to our work, active perception considers motion costs and quality of the perception task. However, the main focus of next best view planning is proposing and selecting views from which a scene can be efficiently observed while in this paper we are interested in computing trajectories from a start towards a predefined goal while striking a good balance between the robot perception performance and most cost. In addition, one of the most important goals of active perception is active localization, which seeks to compute control actions and trajectories that minimize the pose estimation uncertainty. In contrast, our work focuses more on planning for simultaneous perception and motion in a goal-reaching task.
III-D Implicit Neural Representations/Neural Surrogate Model
A surrogate model is an engineering approach employed when directly measuring or computing an outcome of interest is either infeasible or prohibitively expensive. In such cases, an approximate model of that outcome is used instead [64]. With the rise of deep learning, neural or deep learning-based surrogate models have been applied in numerous domains, including earth science [65], image processing [66], fluid dynamics [67], and power grids [68]. In robotics and perception, recent advances in machine learning have produced effective implicit neural representations of spatial information, such as Neural Radiance Fields (NeRFs) [69] and Signed Distance Fields [70]. These representations have been used for tasks ranging from learning multi-object dynamics [71] and defining manipulation planning constraints [72], to enabling reactive robot manipulation [73] and vision-based navigation [74]. They offer compact, efficient storage [74], a continuous geometric representation [72, 74, 73], and can be learned directly from sensor data [75], making them particularly appealing for planning. Moreover, recent years have seen neural representation models trained on broad datasets demonstrate impressive potential for generalizable manipulation [76, 77, 78, 79]. For instance, neural object descriptors (NODs) [77] have emerged as a powerful method for extracting dense, part-level features that generalize across object instances. Building on these directions, our work trains a neural model to map robot configurations directly to the desired perception metric (e.g., object detection confidence score) as it is infeasible and very expensive to compute the perception score for different robot configurations in real world.
Figure 2: Illustration of the PS-PRM pipeline for perception-aware motion planning. Bottom: the environment with multiple robot configurations. Middle blocks: PRM generates candidate paths, a neural surrogate model and occlusion checking run in parallel to evaluate perception scores, and PS-PRM selects the optimal path based on predicted perception quality. Top: the camera view confirms differing perception scores across configurations.
IV Overview of PS-PRM
We propose PS-PRM, a GPU-parallelized roadmap-based motion planner that explicitly incorporates perception quality into planning for high-d o f robots. PS-PRM is designed to operate in both static and dynamic environments, enabling robots to maintain high-quality perception of task-relevant objects or humans while performing navigation or manipulation tasks.
Fig. 2 illustrates the overall PS-PRM pipeline, including roadmap construction, perception score estimation, and path selection based on combined motion and perception costs. At a high level, PS-PRM constructs a PRM over a manifold defined by object-centered visibility constraints, such as line of sight or camera frustum alignment. Each configuration in the roadmap is scored based on its estimated perception quality, using a learned neural surrogate model. To handle cluttered scenes and inter-object occlusions, we further refine the perception score using a ray-casting-based occlusion checking module. Finally, PS-PRM selects the path that maximizes perception quality while minimizing motion cost. We choose PRM over tree-based planners in this work because it simplifies the underlying optimization problem and allows for more efficient node sampling and evaluation in batch, which can be efficiently parallelized on the GPU.
In dynamic environments, where objects or people may move, PS-PRM supports rapid replanning through a GPU-accelerated pipeline that updates perception scores, re-evaluates roadmap validity, and generates new plans in real time. The system is designed to operate efficiently at scale by batching key operations such as forward kinematics, surrogate model inference, and collision checking in environments where radical changes are not expected.
The following sections describe the three main components of PS-PRM in detail. Sec. IV-A introduces the neural surrogate model used to estimate perception scores. Sec. IV-B describes the construction of the perception-score-guided planner. Sec. IV-C presents the GPU-parallelized replanning pipeline.
Figure 3: Illustration of the camera model, depicting both the line of sight (orange) and the camera frustum (purple). A cup is included to demonstrate when it intersecting with the line of sight or falling within the camera frustum.
Figure 4: Pipeline for perception-aware evaluation using a neural surrogate model and occlusion-aware ray tracing. From left to right: the environment contains the target and an occluding object; a surrogate model predicts a dense perception score field; ray tracing incorporates occlusion to refine the score; and camera views at viewpoints A and B illustrate visibility differences. High perception scores correspond to views with minimal occlusion.
IV-A Neural Surrogate Perception Score Models
In real-world environments, it is infeasible to directly evaluate perception scores for all robot configurations due to the time and effort required to physically reposition the robot and collect sensor data at scale. As a result, simulation offers a practical alternative: scoring each configuration by rendering the robot’s camera view and passing it through a perception model. However, this simulation-based approach is still computationally intensive. Each evaluation involves placing the robot in the simulated scene, generating a photorealistic rendering, and running the perception model to obtain a confidence score. As demonstrated in our experiments, rendering high-quality images is time-consuming, especially when evaluating a large number of robot configurations. Moreover, perception scores are environment-dependent; any change in the object’s position requires re-scoring previously scored configurations.
To address these challenges, we propose a neural surrogate model inspired by [80, 81] that approximates perception scores without requiring explicit rendering. Alg. 2 outlines the full pipeline. Given a robot configuration and the target object label, the surrogate model predicts the expected perception performance (e.g.,, detection confidence for a cup). We first apply FK to map the configuration from $\mathcal{Q}$ to the robot’s camera pose. The neural model is then trained to map these camera transforms to perception scores obtained from labeled, photorealistic simulations. Specifically, we generate a dataset by rendering views of objects or humans from varied camera poses in simulation and evaluating each image with a perception model. The surrogate is implemented as a multilayer perceptron (MLP) [82] that takes the relative pose between camera and object, along with an object encoding, and outputs the corresponding perception score. Trained on this dataset, the surrogate enables fast and scalable evaluation during planning. Importantly, because the input is based on camera pose rather than robot-specific joint values, the model generalizes across robots with similar perception setups.
To address object inter-occlusion in cluttered indoor environments, we build upon techniques similar to those in [83, 84] by performing ray-casting towards surrounding objects. This step determines whether the object of interest is occluded by checking for intersections between each cast ray and the Axis-Aligned Bounding Boxes (AABB) of nearby objects. For a new robot configuration, we obtain an occlusion score in addition to the perception score estimated by the neural surrogate model. This occlusion score is computed as the ratio of rays that ultimately do not intersect the bounding box of the target object—–an indicator that something else is blocking the view. If the occlusion score exceeds a chosen threshold, we set the perception score of the neural surrogate model to zero for that configuration. Fig. 4 shows a detailed overview of this occlusion checking step.
Fig. 5 illustrates the overall neural surrogate perception score model. This two-tier approach ensures that the neural surrogate model also accounts for occlusion in an efficient, fast manner, since ray-casting against bounding boxes can be highly parallelized. In practice, relying on bounding boxes provides a reasonable approximation of object geometry, though it may not perfectly capture complex shapes or partial occlusions. Nevertheless, for many motion-planning tasks in cluttered environments, it strikes a good balance between speed and accuracy.
Figure 5: Pipeline for evaluating perception scores for multiple robot configurations in parallel. For each batch of configurations, camera poses are computed using FK, while occlusion checking via ray casting and perception score estimation via a neural surrogate model are performed concurrently. The outputs are combined to yield final perception scores for each configuration.
IV-B Perception Score-Guided Planner
To enable robots to maintain a high score for a given perception task throughout their motion, the PS-PRM planning process consists of two key stages: (1) sampling configurations that satisfy visibility constraints, and (2) constructing, scoring, and planning over a PRM built from these configurations. The first stage ensures that each sampled configuration allows the robot to keep the target object or person within view. The second stage integrates perception scores into the roadmap and performs graph-based planning to find a path that balances motion efficiency and perception quality. We describe each stage in detail below.
IV-B1 Sampling with Visibility Constraints
To maximize perception scores along the planned motion, the robot’s onboard camera must continuously keep the object or person of interest in sight. In this work, we consider two commonly used visibility constraints for maintaining such monitoring, as illustrated in Fig. 3: line of sight and camera frustum. In the line-of-sight model, the robot’s optical axis is aligned with the centroid of the target, effectively keeping the target centered on the image plane (e.g., using a pinhole camera model) [15, 11]. In the camera frustum model, the objective is to keep the target within the camera’s viewing volume, represented as a rectangular pyramid or frustum [85]. These visibility constraints define a manifold within the robot’s free configuration space consisting of configurations where the target remains in view.
More formally, we define the set of all robot configurations that enable the robot to monitor the object or person of interest to lie in a manifold $M$ implicitly represented as the zero-level set defined by $h_{M}(q)=0$ where $q\in\mbox{$\mathcal{Q}$}$, $h_{M}(q):\mathbb{R}^{k}\rightarrow\mathbb{R}^{l}$, and $l$ is the intrinsic dimensionality of the constraint ($l\leq k$). For the line-of-sight model, $h_{M}(q)$ is the distance between the centroid of the object and the line of sight from the robot’s camera. For the camera frustum model, we model $h_{M}(q)$ as $\max(0,\penalty 10000\ \text{sd}_{f}(q,p))$ where $\text{sd}_{f}$ is the signed distance between the object $p$ and the surface of the robot’s camera frustum when the robot is at configuration $q$.
For high-d o f robots, computing motion that considers the aforementioned object-centered models can be achieved using sampling-based motion planning (SBMP) [86] with manifold constraints [87] where planning happens on a subspace of the free configuration space defined by a manifold. Specifically, in the construction of our proposed planner, As shown in Alg. 1, ps-prm first constructs a PRM with states and transitions that satisfy the manifold constraints, using either line-ofsight or camera frustum. The neighbor nodes in the PRM are constrained to lie on the same manifold and are selected based on their geodesic proximity within a certain distance. To sample from the manifold, ps-prm first selects a robot configuration from $\mathcal{Q}_{free}$ and then applies the projection operator to map it onto the surface of $M$. This involves iteratively retracting the point towards a minimum of the constraint function and solving a linear system of equations at each step. Configurations that cannot be projected into valid in-manifold configurations are discarded, and the algorithm resamples until a valid configuration is found or a timeout is reached. This process ensures that the object is placed either within the line of sight or inside the camera frustum.
While enforcing visibility constraints helps keep the target object or person within the robot’s sight, this alone is not sufficient for many real-world perception tasks. Our experiments show that visibility-based approximations often fall short in practice, especially in complex environments where factors such as object shape, occlusions, background clutter, and lighting conditions significantly affect perception performance [1, 2, 3]. To address these limitations, our proposed planner also incorporates perception scores to guide planning toward configurations that support more reliable perception.
IV-B2 PRM construction and scoring
While the visibility constraints provide a bound on necessary conditions for the perception task, they do not consider the actual performance of the task. To remedy this, ps-prm utilizes a scored probabilistic roadmap based on the Probabilistic Roadmap planner (PRM) [40], which incorporates perception scores into planning.
After ps-prm builds the manifold-constrained PRM, the edges initially encode only the distance between configuration nodes under a given metric. The motion cost is also normalized to facilitate its later combination with the perception cost. ps-prm then assigns perception scores to the edges to account for the perception task. Since the two configurations connected by an edge may still differ significantly in distance or perception score, ps-prm also uses intermediate configurations to compute a normalized perception score for an edge, $c_{p}(q_{u},q_{v})\in[0,1]$. ps-prm performs geodesic interpolation between the $q_{u}$ and $q_{v}$ to generate $N$ valid intermediate configurations, $q_{1},q_{2},\ldots,q_{N}$. It then captures the images corresponding to these intermediate configurations from a simulator and computes their normalized perception score $c_{p}(q_{1}),c_{p}(q_{2}),\ldots,c_{p}(q_{N})$. Next, ps-prm calculates the average normalized perception score of the two configurations connected by an edge and their intermediate configurations to obtain the normalized perception score associated with the edge $c_{p}(q_{u},q_{v})$:
$c_{p}(q_{u},q_{v})=\frac{c_{p}(q_{u})+c_{p}(q_{v})+\sum_{i=1}^{N}c_{p}(q_{i})}{N+2}$ (1)
Finally, ps-prm linearly combines the normalized perception score and normalized motion cost for an edge to determine a new edge weight. While other nonlinear combinations are explored, the linear formulation offered a good balance between interpretability and empirical performance. Since the perception score is intended to be maximized while the motion cost is intended to be minimized, we subtract the normalized perception score and add a constant $1$ to ensure that the edge weight remains non-negative, within the range $[0,2]$. The combination is defined as follows, where $\alpha\in[0,1]$ is a factor that balances the perception score and motion cost:
$c(q_{u},q_{v})=1-\alpha\cdot c_{p}(q_{u},q_{v})+(1-\alpha)\cdot c_{m}(q_{u},q_{v})$ (2)
After the PRM is constructed and scored, ps-prm connects the start and goal to the PRM, and uses A* [88] with an admissible heuristic to find a path between the start and goal configurations. In this paper, we demonstrate experiments on object detection and tracking, as well as human face detection, though the approach can be generalized to other perception tasks.
1:Input: Environment $\mathcal{E}$, Robot Model $\mathcal{R}$, Object of Interest $\mathcal{O}$
2:Output: Roadmap $\mathcal{G}=(V,E)$ with perception-aware edge weights
3:Initialize empty graph $\mathcal{G}=(V,E)$
[…]
5:for $i=1$ to NN do
6: Sample $q_{i}$ in free space $\mathcal{C}_{\text{free}}$
7: Project $q_{i}$ onto the manifold using constraint projection
8:if $h_{M}(q_{i})$ is less than a threshold then
9:  Add $q_{i}$ to $V$
10:end if
11:end for
12:for each $q_{u}\in V$ do
13:for each $q_{v}$ within geodesic distance $\delta$ do
14:  Normalize the original distance-based edge weight between $q_{u}$ and $q_{v}$ to obtain the motion cost $c_{m}(q_{u},q_{v})$.
15:  Generate NN intermediate configurations $q_{1},\dots,q_{N}$
16:  Compute normalized perception score $c_{p}(q_{u},q_{v})$ using Eq. (1)
17:  Compute edge weight $c(q_{u},q_{v})$ using Eq. (2)
18:  Add edge $(q_{u},q_{v})$ to $E$ with weight $c(q_{u},q_{v})$
19:end for
20:end for
21:return $\mathcal{G}$
Algorithm 1 Perception Score-Guided PRM Construction (PS-PRM)
1:Input: Robot Configuration $q$, Object $\mathcal{O}$
2:Output: Perception Score $c_{p}(q)$
3:Compute camera transform $T_{c}$ using forward kinematics
4:Query neural surrogate model: $c_{p}(q)\leftarrow\text{NN}(T_{c},\mathcal{O})$
5:Perform ray-casting to check occlusions
6:if occlusion score $\tau$ then
7: Set $c_{p}(q)\leftarrow 0$
8:end if
9:return $c_{p}(q)$
Algorithm 2 Neural Surrogate Model for Perception Score Estimation
IV-C GPU-Parallelized Replanning Pipeline
The proposed perception scoring methods are highly parallelizable. Since the perception scores are computed using a neural surrogate model specifically optimized for GPU inference, we are motivated to implement the entire planning pipeline on the GPU. This GPU-accelerated pipeline performs collision checking, constraint projection, FK, and model inference entirely in parallel, significantly improving runtime efficiency. PS-PRM can therefore operate effectively in dynamic environments where the object of interest may move during execution. In such cases, the planner updates node perception scores in real time and replans the path as needed to adapt to the new configuration. Moreover, communication between the GPU-equipped machine and the robot is lightweight, as only the planned trajectories are transmitted. Fig. 5 shows a detailed overview of the GPU-accelerated process.
When the object of interest moves within the environment, PS-PRM needs to update the perception scores of its nodes and potentially replan the path if necessary. To enable fast, end-to-end replanning, we create a GPU-parallelized pipeline that performs collision checking, constraint projection, forward kinematics, and neural surrogate model inference in batches on the GPU. As the robot executes its planned trajectory, it continuously monitors whether the object has moved. If movement is detected (e.g., by the onboard camera), PS-PRM invokes the parallelized pipeline to re-project the existing nodes and update their perception scores. This involves batch collision checking against the new environment state, constraint projection for any nodes violating newly introduced constraints, batched forward kinematics to update camera transforms, and neural surrogate model inference to recalculate the perception scores.
Nodes that cannot be projected into valid configurations or that collide in the new environment are pruned from the roadmap along with their associated edges. After pruning and re-scoring, the planner re-evaluates the current trajectory by computing a new average perception score; if that score falls below a specified threshold, it triggers a replanning step to find a more suitable path from the robot’s current position to the goal region. If the remaining nodes fail to produce a safe path to the goal, PS-PRM partially rebuilds the roadmap by sampling additional nodes while retaining any valid ones, thus avoiding a complete reconstruction. This leverages each step based on their computational needs, while building a PRM can be performed in seconds, re-evaluating nodes’ collisions and scores can be done at higher frequencies of several times a second. Through this GPU-accelerated parallelization, PS-PRM rapidly adapts to dynamic changes in the environment and maintains both collision-free and perception-optimized trajectories with reduced computation time. This results in a significant reduction in overall computation time and energy usage, enabling deployment in dynamic environments as shown in Experimental Results section Sec. V.
V Experimental Results
In this section, we evaluate our proposed methods in simulation and real-robot experiments using the Hello Robot’s Stretch $2$ [89], Fetch robot, and a UR5 robot mounted on a differential-drive base. We use PyBullet [90] for collision checking, ompl’s Python bindings [91] for planning, and Isaac Sim [92] for simulation and visualization. All experiments were conducted on an Intel i7-12700K CPU and an NVIDIA GeForce RTX4090 GPU. We consider both object detection/tracking task and human face detection task. For the robot’s camera parameters in simulation experiments, we use a resolution of 640 $\times$ 480 pixels, a field of view of 42.5 degrees, and a clipping range of 0.3 meters for the near plane and 6 meters for the far plane.
Motion Planning Object Tracking
Method Preprocess Time (s) Plan. Time (s) Total Time (s) Path Len. (rad) Succ. Rate Det. Rate Conf. Score Track Rate
Post - 0.051 0.051 6.277 99.7% 20.9% 0.531 57.9%
Rejection - 1.264 1.264 10.379 98.9% 29.7% 0.599 74.0%
Manifold-frustum - 0.495 0.495 11.141 98.8% 41.0% 0.597 80.9%
Manifold-line - 2.859 2.859 7.478 90.1% 48.9% 0.482 67.4%
ps-prm-frustum 291.7 0.014 291.714 16.688 96.4% 73.7% 0.761 89.8%
ps-prm-line 341.5 0.018 341.518 15.007 95.7% 80.1% 0.792 91.0%
Neural-ps-prm-frustum 32.08 0.017 32.097 15.814 97.1% 72.8% 0.759 89.8%
Neural-ps-prm-line 31.79 0.021 31.811 13.518 96.3% 79.6% 0.771 90.4%
TABLE I: Planning and perception results for the Stretch robot in a simulated home environment. The results are averaged over 2,500 different motion planning problems. “Preprocess Time” is the time spent (e.g., building the PRM and evaluating the perception scores) before planning. “Plan. Time” is the average planning time. “Total Time” is the overall time including preprocessing and planning. “Path Len.” is the average length of the path. “Succ. Rate” is motion planning’s success rate. “Det. Rate” is the object detection rate. “Conf. Score” is the average confidence score of detections. “Track Rate” is the percentage of the frames that the object is tracked.
V-A Evaluated Methods
We implement four baseline methods as follows:
  • Baseline 1 Post: This baseline method initially performs standard motion planning without accounting for any constraints, and then post-processes the robot’s free joints (e.g., camera pan and tilt) along the resulting path to align the robot’s line of sight with the centroid of the object or person as described in details in Sec. IV-B1. This is achieved through a simple geometric analysis given the position and orientation of the robot at each configuration with respect to the monitoring object or person.
  • Baseline 2 Rejection: This baseline method extends a standard sampling-based planner by modifying its feasibility function to ensure that sampled configurations remain within the robot’s camera frustum.
  • Baseline 3 Manifold-line: This baseline method is an implementation of a manifold constrained motion planner using the distance between the centroid of the object and the line of sight from the robot’s camera’s as $h_{M}(q)$ as described in details in Sec. IV-B1.
  • Baseline 4 Manifold-frustum: This baseline method is an implementation of the manifold constrained motion planner described in Sec. IV-B1 using $h_{M}(q)=\max(0,\text{sd}_{f}(q,p))$ where $\text{sd}_{f}$ is the signed distance between the object $p$ and the surface of the robot’s camera frustum when the robot is at configuration $q$.
We implement four versions of our approach ps-prm, as described in Sec. IV, varying along two key dimensions: the perception model (line-of-sight vs. camera frustum) and whether a neural surrogate is used for scoring.
  • PS-PRM-line: Uses a line-of-sight model to estimate perception quality, with direct evaluation (no neural surrogate).
  • PS-PRM-frustum: Uses a camera frustum model to estimate perception quality, with direct evaluation (no neural surrogate).
  • Neural-PS-PRM-line: Uses a line-of-sight model, with perception quality predicted by a neural surrogate model.
  • Neural-PS-PRM-frustum: Uses a camera frustum model, with perception quality predicted by a neural surrogate model.
For both PS-PRM-line and PS-PRM-frustum, we select $N=5$ and $\alpha=0.75$ in our experiments to strike a balance between action cost and perception performance. It is important to note that our planner is not highly sensitive to these parameters as long as they remain within a reasonable range. Through experimentation, we found that small variations in $N$ and $\alpha$ do not significantly impact the overall performance.
For both Neural-PS-PRM-line and Neural-PS-PRM-frustum, we set the number of samples to $N=5$ and the weight parameter to $\alpha=0.8$ in our experiments, aiming to balance the trade-off between action cost and perception performance. To support this framework, we train a neural surrogate model that predicts perception scores for the relevant tasks.
Specifically, in our experiments, we use the object detection confidence score as the perception score for object detection and tracking tasks, as described in Sec. V-B, Sec. V-E, and Sec. V-F. For human detection tasks, such as those discussed in Sec. V-C and Sec. V-D, we use the confidence score of the human face detector RetinaFace [93] from the deepface [94] as the perception score.
The neural surrogate model is implemented as a Multilayer Perceptron (MLP) [82], consisting of a shared architecture with five fully connected layers, each with 256 hidden units and ReLU (Rectified Linear Unit) activation. The input to the network includes the relative pose (comprising both position and orientation) between the camera and the object, as well as an object encoding that indicates which object the robot is monitoring. We use a one-hot encoding to represent the object being monitored. The network outputs the predicted perception score corresponding to that camera transform and object pose.
This model is trained on a dataset composed of a diverse set of commonly encountered objects and humans from the COCO dataset [95], enabling generalization across a wide range of realistic detection scenarios. We generate a dataset of perception scores as the objects or human are viewed from different camera transform in Issac Sim. Specifically, we sample a set of 1000 camera transforms where the object or human are in the field of view uniformly at random. For each camera transform, we render the images and evaluate the perceptions scores using a perception model with the relevant perception tasks. Then, we train our model on the collected dataset for 200 epochs with an Adam [96] optimizer, learning rate of $1\times 10^{-4}$, and batch size of 512.
V-B Benchmark Results
For simulation experiments, we create an environment using objects commonly found in a home such as a table and a cabinet (shown in Fig. 6). Additionally, we introduce five different objects (a cup, a suitcase, a potted plant, a fork, and a TV) common in a home for the robot to detect and track, further mimicking real-world tasks.
We generate $2,500$ motion planning problems by randomly sampling collision-free start and goal configurations for the base, while adjusting the pan and tilt joints so that the robot is directly facing the tracking object. Specifically, we sample 500 problems for each of five objects, with five variations in the object’s poses within the environment for each set.
Table I presents results across 2,500 planning problems, evaluating metrics such as planning time, path length, success rate. We use YOLOv9 [97] and DeepSORT [98] for real-time perception performance assessment. It is measured by per-frame object detection rate, average confidence score of object detection, and tracking rate.
For the perception task of object detection and tracking, all methods except PS-PRM-frustum, PS-PRM-line, Neural-PS-PRM-frustum, and Neural-PS-PRM-line exhibit detection rates below 50% and confidence scores under 0.6, as they do not directly integrate the tasks into planning. This highlights the importance of our method in directly integrating perception scores into the planning process. In contrast, both PS-PRM-frustum, PS-PRM-line, Neural-PS-PRM-frustum, and Neural-PS-PRM-line achieve an object detection rate above 70%, a confidence score exceeding 0.75, and a tracking rate around 90%, with PS-PRM-line delivering the best object tracking performance among all methods.
It should be noted that constructing the constrained PRM with perception-based edge scores, using approximately 3,000 nodes across different monitored objects and scenes (in this experiment, 5 different objects and 5 locations each), takes on average 341.5 seconds for PS-PRM-line and 291.7 seconds for PS-PRM-frustum. Once the PRM is built, subsequent planning is very fast, as demonstrated in Table I.
The preprocessing times for PS-PRM-frustum and PS-PRM-line are relatively long due to the high computational cost of rendering the scene to compute perception scores. In contrast, Neural-PS-PRM-frustum and Neural-PS-PRM-line achieve significantly shorter preprocessing times by using a lightweight neural surrogate model to directly estimate perception scores, thereby avoiding costly rendering and computationally intensive perception models. This approach is also advantageous in dynamic scenarios, as perception scores can be efficiently updated when the monitored object or human moves.
Notably, the performance of Neural-PS-PRM-frustum and Neural-PS-PRM-line remains comparable to that of PS-PRM-frustum and PS-PRM-line. Specifically, Neural-PS-PRM-frustum and Neural-PS-PRM-line are approximately ten times faster than the version without the neural surrogate model, while maintaining comparable performance, achieving detection rates within 1% and confidence scores within 0.01 of the original. Interestingly, the resulting path lengths are also slightly shorter. This suggests that access to ground-truth perception scores may not be essential, and that a well-trained surrogate model can provide sufficiently accurate estimates to achieve a comparable or even better balance between perception quality and motion efficiency.
Fig. 6 illustrates an example of robot trajectories and corresponding camera view images for three different methods: post, Manifold-frustum, and PRM-perception. Both the robot paths and camera snapshots are sampled uniformly along the full trajectory. The figure demonstrates that the method which directly integrates perception scores, PS-PRM-line, achieves significantly better object detection and tracking performance compared to the baseline methods. Notably, it selects a trajectory that stays slightly farther from the monitored object, resulting in a better viewing angle at the cost of a slightly longer path.
These results demonstrate that integrating perception scores into motion planning, especially through a neural surrogate model, enables efficient planning while maintaining high perception quality. This provides a scalable and effective solution for real-time perception-aware robot motion.
In summary, these experiments highlight that: (a) the proposed approach, PS-PRM, significantly outperforms all baseline methods (b) baseline methods based on only visibility constraints are insufficient for practical perception tasks such as object detection and tracking.
Figure 6: This figure shows one example of the robot trajectory and corresponding camera view images for three different methods (post, Manifold-frustum, and PRM-perception) for the simulation experiment in Sec. V-B. Both the robot traces and camera images are sampled evenly along the robot’s complete path. The robot’s initial position is shown in full opacity, gradually fading as it moves towards the goal. The camera view images are displayed in sequential order from top left to bottom right. If a snapshot does not display a score, it indicates that the object was not detected, and the score is zero.
V-C Human Detection in a Nursing Setting
As shown in Fig. 7, we create a nursing environment with two nurses with masks and one patient. The robot is required to plan a path from the start to the goal position while aiming to continuously detect some key features of the nurses. In this experiment, we consider the face detection task, and we use deepface and RetinaFace for real-time face detection performance assessment. For this multi-person task, we use the midpoint between the two individuals as the monitoring point and the average of their face detection scores to construct the PRM and assess performance. In this experiment, we use the Fetch robot and plan the motion of its base, torso, and camera to perform the assigned task.
For the specific scenario depicted in Fig. 7, we evaluate both Manifold-line and Neural-PS-PRM-line over 100 independent runs. Each run of Manifold-line is given a maximum of 60 seconds to compute a solution. For Neural-PS-PRM-line, we construct a PRM and score its nodes and edges by incorporating the face detection confidence score into the planning process with a trained neural surrogate model.
Table II summarizes the average performance over all runs, reporting metrics such as planning time, path length, face detection rate (for detecting at least one nurse or both nurses), and average face detection confidence score. We observe that Manifold-line tends to choose a shorter path by moving behind the nurses, as shown in Fig. 7, which often leads to poor visibility of their faces. In contrast, Neural-PS-PRM-line consistently favors slightly longer and cluttered paths that provide better visibility of the nurses’ faces.
As shown in Table II, Neural-PS-PRM-line significantly outperforms Manifold-line in both face detection rate and average confidence score. This improvement comes at the cost of increased path length and planning time. Specifically, Neural-PS-PRM-line achieves over 98% success in detecting at least one nurse and over 63% for detecting both nurses, while the baseline method fails entirely in the two-face detection task. An example of the routes selected by our method and the baseline is shown in Fig. 7.
In summary, these experiments confirm that our approach PS-PRM is effective across different environments and is a suitable candidate for planning simultaneous perception and motion for high-DoF robots in practical human-robot interaction applications.
Figure 7: In this simulated nursing environment, the robot plans a path from start to goal while continuously tracking both nurses with masks. The baseline approach Manifold-line selects a shorter path (orange) with poor nurse visibility, while the proposed method Neural-PS-PRM-line takes a longer path (blue) for better detection performance.
Motion Planning Human Detection
Method Plan. Path Det. Rate Det. Rate Conf.
Time (s) Len. (rad) (1 face) (2 faces) Score
Manifold-line 5.31 6.89 31.53% 0.00% 0.49
Neural-PS-PRM-line 32.91 11.51 98.40% 63.29% 0.81
TABLE II: Comparison of Manifold-line and Neural-PS-PRM-line methods in a simulated nursing environment. “Plan. Time” refers to the average planning time, with Neural-PS-PRM-line also accounting for the time required to construct the PRM. “Path Len.” represents the average path length. “Det. Rate (1 face)” and “Det. Rate (2 faces)” indicate the face detection rates for one or both nurses, respectively. “Conf. Score” refers to the average confidence score of face detections for the two nurses.
V-D Planning in a Dynamic Environment
As shown in Fig. 8, we construct a dynamic environment containing a human, a cabinet, a shelf, and other common household objects such as a sofa, table, and chair. The robot is tasked with planning a path from its initial position to reach one of two goal locations (the cabinet or the shelf), while continuously detecting key facial features of the human. This experiment focuses on a face detection and tracking task, utilizing DeepFace and RetinaFace to assess real-time face detection performance. All trials are conducted using the Stretch 2 robot. During execution, the human moves and rotates from side to side, requiring the robot to dynamically select which goal to pursue in order to maintain high face detection performance.
For the specific scenario illustrated in Fig. 8, we evaluate both Manifold-line and Neural-PS-PRM-line over 100 independent runs. Each Manifold-line trial is given a maximum of 60 seconds to compute a plan. In contrast, Neural-PS-PRM-line constructs a PRM and assigns perception-aware scores using a trained neural surrogate model based on face detection confidence. Unlike Manifold-line, which does not account for the human’s dynamic movements, Neural-PS-PRM-line supports online replanning during execution as described in Sec. IV-C. Table III presents the average results across all runs, evaluating metrics such as planning time, path length, face detection rate, and average confidence score.
As expected, Manifold-line consistently selects a single goal and results in very low face detection performance. In contrast, Neural-PS-PRM-line initially chooses the path toward the shelf on the left, as the human begins facing that direction. During execution, the robot replans in real time and switches to the path toward the cabinet on the right as the human turns, thereby maintaining better visibility of the face. As shown in Table III, Neural-PS-PRM-line significantly outperforms Manifold-line in both face detection rate and average confidence score, while also demonstrating successful online replanning. An example of the routes selected by our method is shown in Fig. 8.
In summary, these experiments confirm that our approach PS-PRM is effective for planning simultaneous perception and motion for high-DoF robots in practical human-robot interaction applications in dynamic environments.
Figure 8: Example scenario of planning in a dynamic environment for human face detection. The environment contains a human, a sofa, a table, a chair, a cabinet (Goal 2), and a shelf (Goal 1). The Stretch 2 robot must reach one of the two goal locations while maintaining high face detection performance. The human rotates during the task, causing changes in facial visibility. Initially, the robot plans toward the shelf (Goal 1, orange path) as the human faces left. As the human rotates right, the robot dynamically replans and switches toward the cabinet (Goal 2, blue path) to preserve better face visibility. Our Neural-PS-PRM-line method supports online replanning to adapt to the human’s movements, achieving much higher face detection rates compared to static planning approaches.
Motion Planning Face Detection
Method Plan. Time (s) #Plans Path Len. (rad) Det. Rate Conf. Score
Manifold-line 0.17 1 5.48 74.13% 0.55
Neural-PS-PRM-line 2.49 2.12 10.53 90.53% 0.73
TABLE III: Comparison of Manifold-line and Neural-PS-PRM-line methods in a dynamic, human-aware navigation scenario. ”Plan. Time” refers to the average time spent planning, with Neural-PS-PRM-line including PRM construction. ”#Plans” indicates the average number of replanning instances during execution. ”Path Len.” is the average path length in radians. ”Det. Rate” is the percentage of frames in which the person’s face was successfully detected. ”Conf. Score” represents the average confidence score of detected faces.
Figure 9: In this real robot experiment, the robot navigates from the start to goal while trying to detect a cup on the table. The figure shows the motion paths for two methods: Manifold-line (orange) and PS-PRM-line (blue). Camera views for each method are shown progressively from top left to bottom right.
V-E Mobile Manipulator Planning in Blind Spot Alley
In Fig. 1, we illustrate another scenario where a mobile manipulator (a UR5 robot mounted on a differential-drive base) moves from a start configuration at the right-side corner to a goal configuration at the left-side corner, navigating through a cluttered environment. During this motion, the robot must maintain high perception performance, measured by its ability to continuously track a monitor placed at the far end of the cross lane. We use Neural-PS-PRM-frustum in this case to plan full-body motion for the robot, ensuring that the monitor remains consistently detected with high confidence scores. As shown in Fig. 1, the robot continuously adjusts its arm to maintain a clear view of the monitor with its wrist-mounted camera, while simultaneously navigating its differential-drive base to avoid obstacles. The accompanying camera views confirm that the monitor is successfully detected throughout the entire motion. This capability is broadly applicable, enabling the robot to perceive blind spots in the environment to avoid collisions or continue tracking critical objects or humans.
Figure 10: In this real-robot experiment, the environment includes a nurse, her work tables and platform, a chair, a garbage can, and two cabinets (designated as Goal 1 and Goal 2). The Stretch 2 robot must navigate to one of the two goal locations while maintaining high face detection performance. During the task, the human rotates, causing variations in facial visibility. Initially, the robot plans a path to the left cabinet (Goal 1, orange path) as the human faces left. As the human turns right, the robot dynamically replans and switches to the right cabinet (Goal 2, blue path) to preserve better face visibility. Our Neural-PS-PRM-line method enables online replanning in response to human movement, maintaining strong face detection performance throughout the task.
V-F Real Robot Experiments
V-F1 Static Object Tracking Task
In this experiment, as shown in Fig. 9, we set up an environment where the robot must choose a path, either left or right around the table, to reach the other side. Simultaneously, the robot is tasked with continuously detecting the cup placed on the table. The left path is shorter, but visibility of the cup is poor due to occlusion by the monitor. The right path is longer, as it is partially blocked by the chair, but offers better visibility of the cup.
We compare PS-PRM-line (blue path) with Manifold-line (orange path). As illustrated in Fig. 9, PS-PRM-line takes a longer route to avoid occlusion by the monitor, ensuring continuous detection of the cup, while the Manifold-line opts for the shorter path, which leads to the cup being occluded. This is further supported by the robot’s camera view images (displayed sequentially from top left to bottom right), showing that PS-PRM-line detects the cup in significantly more frames compared to Manifold-line.
V-F2 Dynamic Human Detection Task
In this experiment, illustrated in Fig. 10, we demonstrate the deployment of our system in a realistic nurse training scenario. The environment includes a nurse engaged in her workspace, surrounded by typical elements such as work tables, a chair, a platform, a garbage can, and two cabinets that serve as potential goal locations where the robot needs to grab some tools for the nurse. The task requires the Stretch 2 robot to navigate to one of these locations while maintaining reliable face detection of the nurse As the nurse rotates during the task, the robot must adapt to changes in facial visibility. Our system initially plans a path to the left cabinet (Goal 1) when the nurse faces left, but dynamically replans to the right cabinet (Goal 2) as she turns right, preserving optimal visibility. This experiment highlights how Neural-PS-PRM-line enables perception-aware, online replanning to support continuous coordination between motion ad perception, a critical requirement in nurse training and collaborative medical environments.
In summary, these real-robot experimenta show that PS-PRM is suitable for task requiring simultaneous perception and motion in dynamic environments.
VI Discussion
In this work, we address the problem of planning robot motions for a high-degree-of-freedom (d o f) robot that effectively achieves a given perception task while the robot and the perception target move in a dynamic environment. We propose a GPU-parallelized perception-score-guided probabilistic roadmap planner with a neural surrogate model (ps-prm): a roadmap-based planner that explicitly incorporates the estimated quality of a perception task into motion planning for high-d o f robots. Our method uses a learned model to approximate perception scores and leverages GPU parallelism to enable efficient online replanning in dynamic settings.
[…]
Our case studies highlight the practical value of our method in diverse perception-aware planning tasks: in the UR5 scenario, the arm continuously reorients to maintain a clear view of a monitor despite occlusions; in the Fetch robot experiments, the planner actively selects longer but more perceptually informative paths through nurse-populated corridors; and in dynamic human-robot scenarios with the Stretch robot, the system successfully replans on-the-fly to maintain face visibility despite human movement and rotation.
While our work focuses on object and face detection, the underlying framework is general and applicable to other perception-driven tasks such as semantic mapping, affordance detection, or visual servoing. PS-PRM serves as a bridge between motion generation and high-level perception, enabling robots to make informed decisions that account for both geometry and visibility constraints.
This direction opens several promising avenues for future research. First, while the neural surrogate model significantly reduces computation time, it is currently trained offline under fixed environmental conditions. Future work could explore domain adaptation techniques or continual learning strategies to generalize across environments and sensor configurations. Second, incorporating uncertainty estimates from the surrogate model may enhance robustness, particularly in ambiguous or noisy scenes. Third, learning joint representations that fuse geometry, semantics, and visibility could yield tighter integration between perception and planning, especially in occlusion-heavy settings.
Finally, real-world deployments in assistive and collaborative environments, such as smart homes and healthcare, will require consideration of social and ethical factors. For example, in nursing applications, robots may need to detect faces, interpret gestures, and maintain situational awareness while respecting privacy. Although this work is focused on training simulations, a natural extension is to integrate privacy-preserving constraints into the planning process [99], allowing robots to intelligently navigate and perceive while adhering to user-centered constraints.
In summary, this work demonstrates the effectiveness of perception-guided planning for high-DoF robots in both static and dynamic environments. By jointly reasoning about motion and perception, PS-PRM offers a flexible and generalizable foundation for building more capable, context-aware robotic systems.